A particle A moves in a resisting medium in a straight line such that its distance x from a fixed point O satisfies the equation d^2x/dt^2 + p(dx/dt) + qx = 0, where p and q are constants. Find the condition(s) on p and q such that the motion of A is
(i) simple harmonic.
(ii) damped harmonic.
In the case where the motion is damped harmonic, find
(iii) the damping factor.
(iv) the period of the motion.
1
Expert's answer
2021-01-25T03:10:20-0500

(i) "p = { b \over m} = 0"

Where b is a constant that depends on the medium and the shape of the body


Note: "p" is the damping coefficient


"q= {k \over m}=\omega^2"

where is the elastic constant

(ii)"p = { b \over m}" is not equal to zero


"q= {k \over m}=\omega^2"


(iii) damping ration or damping factot,"\zeta" = "{p \over 2 \sqrt{mk}}"


Note:"2 \sqrt{mk} =" critical damping



(iv) period,T, ="{2\pi \over \omega}"